OpenStudy (anonymous):

simplify: sqrt[3]{x^2y^{3/2}}

OpenStudy (anonymous):

\[\sqrt[3]{x^2Y^(3/2)}\]

OpenStudy (anonymous):

the y is raised to 3/2

OpenStudy (amistre64):

^{...}, not ^(...) for latex

OpenStudy (amistre64):

one property that we need to use is\[\sqrt[n]{...}=(...)^{1/n}\]

OpenStudy (anonymous):

\[\sqrt[3]{x^2y^{3/2}}\]

OpenStudy (amistre64):

another property is:\[(a^p)^q=a^{pq}\]

OpenStudy (amistre64):

yw .... let me know how well you can apply those properties, and ill let you know if its correct :)

OpenStudy (anonymous):

i don't get this at all :(

OpenStudy (amistre64):

\[\sqrt[3]{x^2y^{3/2}}\] lets start by getting rid of that umbrella and changing it into an exponent 1/3, thats the first property i stated \[\Large (x^2y^{3/2})^{1/3}\] now we can use the second property, the exponent of an exponent get multiplied \[\Large x^{2/3}~y^{3/(2*3)}\] now lets reduce the fractions \[\Large x^{2/3}~y^{1/2}\] now depending on what a "simplified" form is spose to look like, they might want you to stick an umbrella over them again .... indexed by the denominator of the exponents \[\Large \sqrt[3]{x^{2}}~\sqrt{y}\]